OpenStudy (anonymous):

Will Give Medal! 6(a+2)+3=3(10-3a) Solve for a

OpenStudy (bioepic):

3

OpenStudy (bioepic):

When you distribute, you get 6a+12+3=30-9a

OpenStudy (bioepic):

6a+15=30-9a

OpenStudy (bioepic):

15a=15

OpenStudy (bioepic):

a=1

OpenStudy (kobeni-chan):

Use the distributive property. On the left side of the equal sign, multiply 6 into a and 2, since they are inside the parentheses. This will make it (6a + 12) + 3 = 3(10-3a) Now distribute/multiply 3 to the values in the parentheses on the right side. This will make it 6a + 15 = 30 - 9a Add 9a to each side. 15a + 15 = 30 Subtract 15 from each side. 15a = 15 a=1

OpenStudy (bioepic):

I did it wrong the first time.

OpenStudy (anonymous):

Thank You!!!

OpenStudy (bioepic):

No problem!